### v1

# CLE-SH: Comprehensive Literal Explanation package for SHapley values by statistical validity

Abstract
Recently, SHapley Additive exPlanations (SHAP) has been widely utilized in various research domains. This is particularly evident in medical applications, where SHAP analysis serves as a crucial tool for identifying biomarkers and assisting in result validation. However, despite its frequent usage, SHAP is often not applied in a manner that maximizes its potential contributions. A review of recent papers employing SHAP reveals that many studies subjectively select a limited number of features as ’important’ and analyze SHAP values by approximately observing plots without assessing statistical significance. Such superficial application may hinder meaningful contributions to the applied fields. To address this, we propose a library package designed to simplify the interpretation of SHAP values. By simply inputting the original data and SHAP values, our library provides: 1) the number of important features to analyze, 2) the pattern of each feature via univariate analysis, and 3) the interaction between features. All information is extracted based on its statistical significance and presented in simple, comprehensible sentences, enabling users of all levels to understand the interpretations. We hope this library fosters a comprehensive understanding of statistically valid SHAP results.
Keywords SHapley Additive exPlanations  $\cdot$ Statistical Validity  $\cdot$ Visualization
1 Introduction
1.1 Background
With the development of machine learning, the interpretation of models, specifically understanding the importance of each feature, has gained significant attention in biomedical areas. In these areas, datasets are often vulnerable to data bias, meaning high performance alone is not enough to validate the use of machine learning[47, 12, 9]. Interpretation that makes sense to experts can help reduce the potential risks of this bias. As machine learning utilizes non-linear and complex interactions between features that conventional regression analysis cannot handle, interpretation is often used to understand the mechanisms of target diseases[21, 44, 54]. Additionally, interpretation is sometimes used as a primary goal to identify biomarkers among the vast numbers of genetic sequences or microbiomes[50, 42, 23, 5].
Interpretation of machine learning models built for tabular datasets can be done using two approaches: model-dependent and model-agnostic. The model-dependent approach includes indices such as permutation importance or the Gini-impurity of tree-ensemble algorithms like XGBoost and LGBM, which are popular these days. However, these measures provide only absolute importance, offering limited information compared to regression analysis. Therefore, the model-agnostic method, represented by SHapley Additive exPlanations (SHAP), is predominantly used for interpretation. SHAP analysis is an interpretable AI method based on game theory, elucidating how each feature affects predictions under various conditions. By providing explanations for each sample, it allows for a deeper understanding of feature importance, feature importance distribution by its feature value, and interaction analysis[30].
Between April 1st, 2024, and April 7th, 2024, by searching for (SHAP[Title/Abstract]) AND (Machine Learning[Title/Abstract]) on the PubMed site, 38 papers that use machine learning and SHAP together were published. We analyze how papers in the biomedical fields utilize SHAP in Figure1 and Appendix Table 1. Among 33 accessible manuscripts, 84.8% of the papers include SHAP summary plots, which summarize the SHAP value distribution for each feature by its rank [2, 15, 52, 14, 18, 28, 41, 25, 45, 53, 1, 46, 48, 40, 20, 31, 39, 19, 43, 17, 49, 33, 26, 34, 38, 8, 24, 10, 7, 51, 35, 16, 11, 32, 6, 27, 29, 37].
Figure 1: Investigation on the SHAP analysis usage in recent papers in medical fields
1.2 Problem Definition
We have identified three major issues with the current application of SHAP in state-of-the-art research.
1) Lack of statistical validation: Validation is critical in any field, especially in medicine. However, the current usage of SHAP in biomedical research lacks statistical validation and is often arbitrary. For example, Figure1.d shows that among papers that did not display all features in the SHAP summary plot, only 15% used feature selection methods. The majority chose arbitrary numbers like 10, 15, or 20 features without justification (Figure1.e). Only one paper applied statistics to SHAP values, using them for Principal Component Analysis rather than to validate SHAP results [39]. Without statistical tests, biomedical experts, particularly those unfamiliar with machine learning, may find it difficult to trust and utilize SHAP results.
2) Complexity of interpretation: While SHAP summary plots and other visualizations may be intuitive for data experts, they can be challenging for medical experts to interpret. The mixed colors in SHAP summary plots or SHAP dependence plots can obscure clear patterns, making it difficult to discern even highly ranked features.
3) Absence of non-linear analysis: SHAP is often used similarly to traditional feature importance methods, focusing only on ranking features. This neglects the unique capability of SHAP to provide interactive analysis through SHAP dependence plots. As shown in Figure1.b, 81.8% of papers did not include SHAP dependence plots. Even among the 18.2% that did, most labeled scatter plots of single features as SHAP dependence plots.
These factors hinder the utility of SHAP analysis in providing interactive analysis and being referenced by medical experts who are not familiar with machine learning. However, there is no clear consensus or set of tools on how to extract statistically valid results from SHAP with comprehensive illustrations. Considering the different categories of feature variables and the varying patterns of SHAP values by feature, it is difficult and cumbersome for each researcher to do this individually.
1.3 Proposed Approach
Figure 2: Schematic of the library
We propose a Python library package that allows any user to easily access and apply proper statistical methods to their SHAP values, generating a comprehensive literal report PDF file that even non-experts in data science can easily understand. Meanwhile, we allow any user to utilize this package as they intend by setting hyperparameters.
1) Statistical tests to analyze SHAP values: In Sections 2 to 5, we introduce a procedure to analyze SHAP values with statistical significance for any type of tabular dataset. First, we define the number of important features to focus on, which involves feature selection by SHAP values (Section 2). Next, we determine the data type of each important feature, whether it is binary, discrete, or continuous (Section 3). Considering the type of each feature, we suggest univariate analysis to analyze each feature individually (Section 4). Finally, we design an interactive analysis to filter out statistically significant results from SHAP dependence plots (Section 5).
2) Summary report generation: In Section 6, we summarize the statistically significant results into a report intended to be comprehensive for non-experts in data analysis. To achieve this, instead of only displaying plots, we also generate literal sentences that explain the significant patterns within the plots.
1.4 How to use the library
The description and installation guide for the CLE-SH package is accessible at: xxx.xxx.xxx. After installation, importing the package with import clesh will enable the analysis by executing clesh(features, SHAP values, label name) as in Figure2. Here, features indicates the dataframe containing the feature dataset used to train/test the machine learning model and calculate SHAP values. SHAP values are SHAP values which match to features. label name is required to generate more comprehensive literal sentences.
1.5 Examples of the application
To validate the generalizability and reproducibility, we test our package using examples from five different datasets from various biomedical areas. These datasets vary in the number of samples, features, feature types, and obtainable performance (Table 1, Appendix 1) [4, 36, 13, 3, 23]. We collect SHAP values by test features in each 5-fold cross-validation, ensuring that every data sample was included in the results. The XGBoost algorithm, a widely used tree-based ensemble method popular in biomedical research, is used for the machine learning model (Appendix 2). We use the area under the ROC curve (AUC) to calculate the performance for each dataset. Details of our testing environment are provided in Appendix 3.
In this study, our purpose is to test the results of our library, rather than to interpret them in terms of biological or clinical knowledge. Therefore, we do not include any analysis of the results from that perspective, and some features are referred to by their feature numbers rather than names. When users actually use this package, if the input data file includes the names of the features, the results will be generated based on those names.
Table 1: Data Characteristics. MS*=Metabolic Syndrome, BC*=Breast Cancer, HF*=Heart Failure, DR*=Diabetic Retinopathy, IBD*=Inflammatory Bowel Disease
Label Name MS* BC* HF* DR* IBD*
# of samples 2009 198 299 1151 1569
# of positive 822 47 96 611 702
Average AUC 0.96 0.72 0.90 0.77 0.99
# of features 17 31 12 19 283
Binary 7 0 5 3 0
Discrete 1 0 1 0 0
Continuous 9 31 6 16 283
2 Number of important features
Figure 3: Graphical explanation of determining the number of important features using SHAP with examples. a-c. Visual explanation of the feature selection process. d. Feature selection plot derived from the BC dataset and e) from the DR dataset.
Displaying and analyzing every feature might be inefficient as not all features are important and it can harm the effective delivery. As shown in Figures 1.d and 1.e, the number of features displayed in SHAP summary plots is usually decided to be 10 or 20 without validation, such as feature selection. Although justifying the number of important features is crucial, it is very cumbersome for analysts to perform feature selection every time. This inconvenience arises particularly because justifying with the best model algorithm might increase the computational burden. As a result, even when feature selection is used to validate the number of important features to analyze, the algorithm is often based on much simpler and smaller models that are not related to SHAP values or the final predictive model used to collect SHAP values. To address this issue, we use a feature selection method oriented around SHAP values [22]. By automating this process, the justification for the number of important features is conveniently provided.
As statistical significance is a key factor in this study, we compare the average absolute SHAP values between features. As shown in Figure3.a, we calculate the statistical significance of difference between adjacently ranked features. When both groups of SHAP values satisfy the normal distribution test by the Shapiro Wilk test, a paired t-test is used (parametric); otherwise, the Wilcoxon rank sum test is used (non-parametric). By cutting the group of features by where statistical significance occurs, we determine the number of important features according to the number of cuts, as shown in Figure3.b. The minimal number of important features (candidate_num_min, default=10) and the maximal number of important features (candidate_num_max, default=20) are set by hyperparameters, as having too many important features can be ineffective and too few can be less informative. Between candidate_num_min and candidate_num_max, we select the number that shows the highest difference from the next cut, as shown in Figure3.c. Statistical significance is determined p-value<0.05, which 0.05 can be changed by adjusting the hyperparameter p_feature_selection. In the clesh_result folder (see Figure2), the Figureshowing the number of important features according to the number of cuts, as in Figures 3.d and 3.e, will be saved.
3 Identification of feature data type
Defining the type of features is necessary to design an appropriate statistical test pipeline. In this paper, we define three different types of features: binary, discrete, and continuous. By calculating the unique values of each feature, the feature type can be easily determined. If the unique feature values are only two, the feature is classified as binary. If the list of unique feature values exceeds a defined threshold (hyperparameter cont_bound), the feature is classified as continuous; otherwise, it is classified as discrete. We set the default for cont_bound to 10. In Table 1, the number of features for each data type is listed.
4 Univariate analysis
Based on the type of each feature by Section 3, we plot figures which can generally describe the distribution of SHAP values, while analyzing the univariate impact with statistical significance. This analysis is done only for important features selected in Section 2, but can be expanded by setting the hyperparameter manual_num. The standard for statistical significance in this section can be controlled by the hyperparameter p_univariate. In Table 2, we summarize the result of univariate analysis for five datasets. More examples like as Figure4 are illustrated in Appendix 4. Any figures drawn in this section are saved in the folder, clesh_results/univariate_analysis as in Figure2.
Table 2: Univariate analysis in each dataset. Linear, Quadratic, and Sigmoid in the first column imply the number of features that fit best to each function with statistical significance.
Label Name MS* BC* HF* DR* IBD*
# of important features 17 11 12 13 12
Binary 7 0 5 0 0
Discrete 1 0 1 0 0
Continuous 9 11 6 13 12
Linear 0 3 0 0 0
Quadratic 2 2 2 3 0
Sigmoid 7 6 4 10 12
4.1 Binary and Discrete Types
We provide box plots of SHAP values in each category, as shown in Figures 4.a and 4.b.
Inside category: To understand an overview of how each category influences the label, SHAP values of each category are tested to determine whether the average is significantly higher or lower than zero. For each group, normality is tested using the Shapiro-Wilk test; if normal, an one sample t-test is applied, and if not, the Wilcoxon signed rank-sum test is used.
Between categories: The distribution of importance in each categories is compared to understand which category shows different pattern from another. For binary features, two sample t-test is applied when both groups are normal by Shapiro Wilk test, and if not, Mann-Whitney U test is applied. For discrete features, one-way ANOVA (parametric) or the Kruskal-Wallis test (non-parametric) is used to identify any statistically significant differences between categories. If a significant difference is found, Tukey’s HSD is used to identify which categories differ significantly ($\textit{p}<0.05$). The result of Tukey’s HSD is represented as in Figure4.c.
4.2 Continuous Types
SHAP values of continuous features are difficult to interpret using conventional statistics, as the relationship can be very complex to interpret or feature values can be too sparse. To balance the amount of information and utility, we first provide scatter plots (SHAP dependence plots without interaction analysis). Then, we apply three different functions for interpretation: linear (Eq. 1), quadratic (Eq. 2), and sigmoid (Eq. 3). By setting feature values as the input variable and SHAP values as the output, each function is fitted, and its statistical significance is checked. As our goal is to determine which function best represents the impact of a feature, we define statistical significance by the p-value of the coefficient of the input (x), denoted as a. If more than one function type is statistically significant, the Root Mean Square Error from real data points is used to select the best function type. The selected function is displayed in the scatter plot. As shown in Table 2, the function type that fits the most is the sigmoid function, followed by quadratic and linear. This indicates that the relationship between feature values and SHAP values is usually non-linear, underscoring the need to analyze more than just the SHAP summary plot. An example of the linear function is illustrated in Figure4.d, the quadratic function in Figure4.e, and the sigmoid functions in Figures 4.f and 4.g.
[…]
Figure 4: Examples in univariate analysis. a. Binary type (Dataset: MS), b. Discrete type (Dataset: HF), c. Result of Tukey-HSD test (Feature from Figure4b), d. Continuous type fitted to a linear function (Dataset: BC), e. Continuous type fitted to a quadratic function (Dataset: DR), f. Continuous type fitted to a sigmoid function (Dataset: MS), g. Continuous type fitted to a sigmoid function (Dataset: IBD). Other examples can be found in Appendix 4
5 Interaction analysis
Interaction analysis is performed for important features in Section 2, namely, target features. To define the interaction feature of each feature, we use the built-in function of SHAP library, shap.utils.approximate_interactions().
Although interaction analysis is a unique strength of SHAP compared to other feature importance methods, defining interactions can be challenging due to the numerous combinations of data types and relationships. We believe that this complexity, along with the inconvenience of manually examining each SHAP dependence plot, limits the potential of SHAP’s interaction analysis (Figures 1.b and 1.c). Rather than overwhelming users with information, we aim to simplify the pipeline and recommend specific interactions to focus on by illustrating the statistical significance of certain patterns. We plot the SHAP values of the target feature separately based on the interaction feature’s value as in Figure5. Any figures drawn in this section are saved in the folder, clesh_results/interaction_analysis. Additionally, we perform a statistical test to determine whether the value of the interaction feature has a significant impact on the SHAP values of the target feature. The standard for statistical significance in this section can be controlled by the hyperparameter p_interaction.
Figure 5: Examples in interaction analysis. a. Box plot of ’Albuminuria’ with the interaction feature ’Triglycerides’ (Dataset: MS). b. Tukey’s HSD result for samples where Triglycerides is lower than the average. c. Tukey’s HSD result for samples where Triglycerides is higher than the average. d. Scatter plot of Waist Circumference with a sigmoid function fitted by univariate analysis (Dataset: MS). e. Scatter plot of Waist Circumference with the interaction feature ’sex’. Other examples can be found in Appendix 5
1. Target: Binary/Discrete, Interaction: Binary/Discrete. In each category of the interaction feature, the SHAP values of the target feature between different categories of the target feature are compared. The statistical test follows the method described in Section 4.1. When a target feature is binary, a two-sample t-test (parametric) or Mann-Whitney U test (non-parametric) is applied. When a target feature is discrete with multiple categories, a one-way ANOVA (parametric) or Kruskal-Wallis test (non-parametric) is used, followed by Tukey’s HSD post-hoc analysis to identify statistically significant differences between categories.
2. Target: Binary/Discrete, Interaction: Continuous. When the interaction feature is continuous, analyzing by its range can be difficult to interpret. To simplify, we group interaction values into two categories: those smaller than the average and those larger than the average. This allows us to treat the interaction feature as a binary feature and apply the analysis described in Section 5.1.
In Figure5.a, an example of box plots in interaction analysis is illustrated (dataset: MS). The target feature ’Albuminuria’ shows a negative slope in its SHAP values, regardless of the interaction feature’s value (Triglycerides). However, in Figures 5.b and 5.c, we observe that the SHAP values, where ’Albuminuria’s feature value is 2, show a statistically significant difference from other SHAP values only when Triglycerides is above its average. This indicates that the impact of Albuminuria = 2 is more significant when triglycerides are above the average.
3. Target: Continuous, Interaction: Binary/Discrete. SHAP values of the target feature are grouped by the category of the interaction feature. Each group of SHAP values is then regressed by the target feature values using the function type selected in Section 4.2 (among Eq.1 ,2 or 3). Functions that are fitted with statistical significance will be displayed in the scatter plots as shown in Figure5.d. If only one group (one category of the interaction feature) shows statistical significance for the fitting, there will be no comparison between categories of the interaction feature. If more than one group shows statistical significance for the fitting, a comparison is made to determine whether the pattern depends on the category of the interaction feature.
However, comparing two different function equations is complicated because the difference depends on which range of target feature values is compared. And analyzing the difference by its location complicates the interpretation and deviates from our goal of comprehensiveness. Therefore, we only analyze the overall margin between the fitted functions of each group. For every target feature value, we compare the SHAP value regressed by each group using a paired t-test (parametric) or Wilcoxon rank-sum test (non-parametric).
In Figure5.d, although the scatter plot is fitted into sigmoid function, it is not sufficient to explain the three stages of distribution. Figure5.e shows that SHAP values of waist circumference depend on sex, and the scatter plot in Figure5.d represents the mix of two different sigmoid functions ($\textit{p}<0.05$).
4. Target: Continuous, Interaction: Continuous. As in Section 5.3, samples are divided into two groups based on whether the interaction value is below or above its average. The analysis follows the same procedure as in Section 5.3, treating the interaction feature as binary.
6 Conclusion
To the best of the authors’ knowledge, this study is the first to develop a Python library that automatically performs statistical analysis on SHAP values. As illustrated in Figure1, the library automatically generates a report file summarizing all statistically significant patterns. Report PDF files for each dataset are included within the source code.
Proper recognition of the package’s limitations is essential for correct usage. Real-world features are highly interactive and complex, meaning that SHAP analysis applied to machine learning models cannot illustrate every inherent detail. The package presented in this paper summarizes statistically significant SHAP values, but the implications of these significant patterns should be interpreted by domain experts for each specific dataset. It is important to note that interactions between features are not represented by univariate analysis, and even interaction analysis may not fully elucidate the mechanisms underlying feature interactions. We hope that this paper contributes not by finalizing the biomarker discovery process, but by extracting meaningful insights from machine learning models that can guide more targeted research projects.
For future work, we suggest improvements in visualization and interface design. Additionally, we plan to expand the range of analyses in both univariate and interaction contexts. While the current version of the library does not pose a significant computational burden, as it only analyzes precomputed SHAP values, future enhancements might require consideration of strategies to reduce computational load as the library becomes more complex.